OpenStudy (anonymous):

Factor completely 144x2 – 120x + 25

OpenStudy (anonymous):

\[ a^2 - 2 ab + b^2 =(a-b)^2 \]

OpenStudy (anonymous):

\[ 144x^2 – 120x + 25= (12x)^2 -2 (12)(5) + 5^2=(12x-5)^2 \]

OpenStudy (anonymous):

can you show me the steps so i can learn how to do it

OpenStudy (anonymous):

Do you know the formula \[ (a-b)^2 = a^2 - 2 a b + b^2 \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@eliassaab is the answer (12x – 5)2 thats what i got

OpenStudy (anonymous):

yes